# PropNet: a White-Box and Human-Like Network for Sentence Representation

### v1

@@ 1 Introduction
Sentence representation research has been carried out for over half a century, evolving from regular expressions [20], statistical encoding methods [24; 29; 43; 26] to neural network embeddings [35; 9; 28; 27; 18; 40; 19; 32]. In recent years, embedding models have dominated the research community, and also have achieved significant success in industry applications. However, their black-box nature and reliance on large-scale data-driven training raise concerns about bias [10; 12; 11; 37], trust [41; 42], safety [7; 39] and other issues [8]. Researchers have attempted to improve the interpretability of embedding models [25; 33] and remove their biases [23; 22], but these problems remain fundamentally unsolved.
In the cognitive science community, the question of how our brains represent a sentence has drawn a great deal of research interest. When humans read and comprehend a sentence, they decompose it into propositions. These propositions are simple sentences with no more than one verb. The meanings of these propositions are stored in long-term memory as a hierarchical network, which can be recalled for use in downstream tasks [31; 16]. The brain’s verbal short-term memory possesses a specialized subsystem. This subsystem is responsible for splitting propositions and handling the syntactic features of each proposition as the eyes sweep across the text[13]. Moreover, the cognitive load in sentence comprehension is primarily attributed to syntactic complexity, such as the number of verbs, rather than superficial length, such as the addition of adjectives and other modifiers [21]. This implies proposition is a processing unit for sentence comprehension. Inspired by these findings, in this work, we propose a purely white-box and human-like sentence representation network named PropNet.
@@
PropNet constructs a hierarchical network based on the propositions comprising a sentence. The generation of PropNet involves four phases: splitting, parsing, representing and merging. The first two phases employ a mechanism analogous to short-term memory, whereas the last two phases function akin to long-term memory. Figure 2 provides an example of PropNet.
PropNet is purely white-box since it is completely interpretable and transparent. It is also human-like because it has a hierarchy network structure to save the meaning in a sentence. With respect to semantics, dislike embedding models which focus more on the relationships between words and their contextual information, PropNet focuses on the true happenings in the world as expressed by a sentence. Therefore, given a proposition, eight dimensions related to a verb are considered: Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, Source. Moreover, PropNet keeps the relationships among propositions in a sentence to convey a complete and coherent meaning. The principles behind PropNet’s design are very natural, without specially crafted artificial features. The comparison of PropNet and other representation methods is summarized in Table 1.
@@ 2 Relative Work
Neural network embedding approaches represent words as vectors in a continuous vector space. Representative methods include Word2Vec [35], FastText [9; 28; 27], Glove [38], BERT [18], SBERT [40], SimCSE [19] and AoE [32]. They train neural networks on an extremely large corpus, using different Language Modeling (LM) tasks as the objective. Since embeddings are automatically computed by backpropagation, feature engineering is completely eliminated. In addition, the embeddings provide deep semantic understanding of the text, as they capture the semantic aspects of language due to LM tasks as the training goal.
Embedding methods are considered as black-box algorithms because the embedding vectors are generally unexplainable. This characteristic surreptitiously encodes social biases contained in the training corpus [10; 12; 11; 37], and also raises concerns about trust [41; 42], safety [7; 39] and other issues [8]. Numerous efforts have been made to improve interpretability [25; 33] and remove biases [23; 22], but these problems have not been fundamentally solved.
All the aforementioned methods are not human-like, as their representation formats are either strings or vectors. Table 1 summarizes the characteristics of these methods.
@@ 3 Methods
Table 2: Splitting rules for a Prop2 sentence according to its clause types. The condition token.pos_ in (VERB,AUX) is also required for all rules. We omit it to improve the clarity of the table. The examples follow the format: Input Sentence -> (Main Proposition, Subordinate Proposition). The bold text in an example is the extracted verb token for the subordinate clause. Note that identifier_ccomp, identifier_pcomp, xcomp_identifier and identifier_csubj serve as placeholders for the relevant clauses.
The PropNet model comprises four key components: splitting, parsing, representing, and merging. Specifically, the splitting and parsing processes correspond to activities occurring in short-term memory, whereas the representing and merging processes align with those in long-term memory. Subsequently, we introduce a method for comparing two sentence representations, which generates a difference vector as output.
@@
  • P3+ indicates the set of pairs that includes at least one Prop3.
@@ 3.1 Splitting
For Prop0 and Prop1, splitting is unnecessary. At this stage, the input sentence itself serves as the output proposition. For Prop2, split the input sentence into two Prop1 propositions. Table 2 presents the splitting rules for different clause types. These rules extract the verb token of the subordinate clause. Then a subordinate proposition is generated by iterating over the verb token’s subtree provided by spaCy parsing tool  [^1] For more details, see the spaCy documentation: https://spacy.io/usage/linguistic-features/#pos-tagging. For instance, in the sentence “She thinks this is a good idea”, the splitting rules identify “is” as the verb token for the subordinate clause. The token “is” owns a subtree “this is a good idea” by dependency parsing.
We set clause identifiers to record relations between propositions. In the above example, the main clause becomes “She thinks identifier_ccomp” where identifier_ccomp points to “this is a good idea”. The recorded relations are crucial for merging proposition representations into a unified network.
If the subordinate clause is non-finite, the subject of the main clause is spliced to the subordinate clause. For instance, in “she wants to dance”, the subordinate proposition “she dance” uses the subject of the main proposition “she wants xcomp_identifier”. If the subordinate clause is finite, no subject insertion is required since the clause already contains its own subject. The splitting rules for relative clause are more intricate and are detailed in Appendix A.2.
Situation becomes significantly more complex for Prop3+. We employ a backtracking algorithm to decompose the input sentence into Prop1 propositions by constructing a binary tree, as shown in Figure 1. The splitting method for Prop2 is invoked during each splitting step to extract a subordinate proposition Prop1, serving as the leaf node of the binary tree. Then the leaf node invokes the representation component to construct a hierarchical network, which is explained in Section 3.3.
@@ 3.2 Parsing
This phase extracts key tokens from a proposition with respect to specific dimensions. The intuition of designing these dimensions is that we require they can describe the meaning of a single action that happens in the real world, which is the key to understand the meaning of a complex sentence consisting of a series of actions. The action type in a proposition can be physical, mental or social [44]. Therefore, eight critical dimensions related to the verb of a proposition are considered: Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, Source. The extracted tokens under each dimension are either describing a physical entity like “car” or an abstractive concept like “brand”. Two dimensions Attribute and Part_of, are considered for the entities or concepts extracted for action-related dimensions. Explanations for these dimensions and extracting primary rules  [^2] Because of the variety of language, even a Prop1 sentence can have an extremely flexible structure. Therefore, minor rules are handcrafted for extraction which can be seen in the codes. with examples are listed in Table 3. The prepositions used by primary rules in the table are omitted for brevity. See Appendix A.3 for a detailed list of these prepositions. Extraction is implemented by resorting to spaCy part-of-speech tagging and dependency parsing tools  [^3] For more details, see the spaCy documentation: https://spacy.io/usage/linguistic-features/#pos-tagging.
@@ 3.3 Representing
As brain’s long-term memory stores information as a hierarchical network [16; 6; 30], PropNet represents a proposition using a six-level hierarchical network. Figure 2 (b) provides an example of how the sentence “A young man is playing an instrument in the garden” is represented. The network comprises three types of nodes: evolutionary nodes, developmental nodes and instance nodes which are represented as red, orange and green colors, respectively. In designing the network, we incorporate ideas similar to Nature Design and Nurture Beliefs , which are outlined in [44]. Evolutionary nodes correspond to Nature Design, which reflects human nature and is derived from heredity. They are denoted with a prefix “#”, like #action. Each dimension in the parsing phase has a corresponding evolutionary node. Developmental nodes correspond to Nurture Beliefs, which are learned from experience, and are named with a prefix “_”, such as _young. Each word in a sentence corresponds to an orange developmental node, which is pointed by a green instance node.
Instance nodes play a crucial role in preserving the meaning of a proposition, enabling PropNet to have potentials in addressing coreference resolution and representing contextual information effectively. They have the following three types:
@@
Figure 2 (a) provides a more lucid depiction of the six-level hierarchical structure of PropNet. An action instance node functions as the root and triggers #action. This action instance node is unique because there is only one verb in a proposition. Edges from #action connect evolutionary nodes #subject, #object, #where, #aux_obj, #goal, #source, and #reason, each of which corresponds to a parsing dimension described in Section 3.2. These nodes will establish connections with instance nodes that are created for the tokens parsed from each dimension. At level 4, these instance nodes activate two evolutionary nodes, #attr and #part_of, which align with the parsing dimensions Attribute and Part_of respectively. Similarly, #attr and #part_of establish links with instance nodes created for the tokens parsed under Attribute and Part_of.
@@
If during the parsing phase, no words are associated with a dimension, its corresponding evolutionary node will invoke an instance node that points to a default developmental node, _unknown. In Figure 2 (b), only #subject, #object and #where have valid instance nodes, which point to _man, _instrument and _garden, correspondingly. All other evolutionary nodes pointed to by #action are connected to instance nodes that point to _unknown, which is marked by a purple box in the figure.
@@ 3.5.1 Pair Type P1-
The comparison dimensions are formed by the Cartesian product of {#subject, #object, #where, #aux_obj, #goal, #source, #reason} and {$1$, #attr, #part_of}, further extended by {#action, #action|#subject|#where, #action|#object|#where}. For example, the dimension #action|#subject|#attr represents the attribute of an entity or concept that serves as the subject within the action. In addition, the dimension name reveals the route to locate the corresponding entity or concept instance nodes within PropNet. The explanations of these 24 dimensions are provided in Appendix B.1.
By comparing the developmental nodes indicated by the dimensional instance nodes, we obtain a similarity code for each dimension. Concatenating these codes in sequence produces a difference vector $\mathbf{v_{1}}$. The similarity code primarily takes three values: 0 for identical, 1 for similar, 2 for different. A large language model (LLM) acts as the knowledge base to compute the similarity degree between two words. Details on LLM prompts, additional similarity codes, and examples for each code are provided in Appendix B.3. Extending PropNet to compute lexical similarity is left in future research.
@@ 3.5.3 Pair Type P3+
As a Prop3+ sentence can contain multiple propositions, we estrict our consideration to its first four propositions in the order of appearance. Therefore, the difference vector $\mathbf{v_{3}}$ for P3+ has a size that is twice the length of $\mathbf{v_{2}}$ for P2. We calculate the overlap degree as the number of overlapping words between two propositions. The alignment strategy involves iterating through all the propositions of the first and second sentences, and then choose the pair with the maximum overlap degree. After that, we remove the selected propositions from the list and continue to apply the aligning strategy to the remaining propositions until the maximum overlap degree is zero.
Let’s exemplify the above process, considering the pair: “Three men are jumping off a wall” and “Three young men run, jump, and kick off of a Coke machine”. The second sentence consists of three propositions: “three young men run”, “three young men jump”, and “three young men kick off of a Coke machine”. After the alignment process, its proposition list is reordered to “three young men jump”, “three young men run”, “three young men kick off of a Coke machine”. The first sentence is compared with “three young men jump”, resulting a difference vector $\mathbf{v_{1}}$. Since “three young men run” and “three young men kick off of a Coke machine” have no counterparts, we concatenate $\mathbf{v_{1}}$ with two vectors of the same length as $\mathbf{v_{1}}$, padded with 2. Finally, the resulting vector is padded with 0 until it reaches the required length.

### v3

@@ 1 Introduction
Sentence representation research has been carried out for over half a century, evolving from regular expressions [23], statistical encoding methods [27; 36; 50; 32] to neural network embeddings [42; 10; 35; 34; 20; 47; 22; 39]. In recent years, embedding models have dominated the research community, and also have achieved significant success in industry applications. However, their black-box nature and reliance on large-scale data-driven training raise concerns about bias [11; 14; 13; 44], trust [48; 49], safety [8; 46] and other issues [9]. Researchers have attempted to improve the interpretability of embedding models [30; 40] and remove their biases [26; 25], but these problems remain fundamentally unsolved. Moreover, in the process of training models, massive amounts of manually labeled data have been generated. However, researchers typically focus solely on the training value of these labeled data, neglecting the cognitive values that exist within them. These cognitive values hold significant potential for enhancing human self-understanding.
In the cognitive science community, the question of how our brains represent a sentence has drawn a great deal of research interest [37; 7; 6; 21; 24; 29; 31]. When humans read and comprehend a sentence, they decompose it into propositions. A proposition is a simple sentence with no more than one verb. The meanings of these propositions are stored in long-term memory as a hierarchical network, which can be recalled for use in downstream tasks [38; 18]. The brain’s verbal short-term memory possesses a specialized subsystem. This subsystem is responsible for splitting propositions and handling the syntactic features of each proposition as the eyes sweep across the text [15]. Moreover, the cognitive load in sentence comprehension is primarily attributed to syntactic complexity, such as the number of verbs, rather than superficial length, such as the addition of adjectives and other modifiers [24]. This implies proposition is a processing unit for sentence comprehension. Inspired by these findings, in this work, we propose a purely white-box and human-like sentence representation network named PropNet.
@@
PropNet constructs a hierarchical network based on the propositions comprising a sentence. The generation of PropNet involves four phases: splitting, parsing, representing and merging. The first two phases employ a mechanism analogous to short-term memory, splitting a sentence into propositions and parsing key tokens from each proposition according to specific dimensions. The dimensions include Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, and Source. The last two phases function akin to long-term memory, representing a proposition as a hierarchical network and merging all networks into a unified network to express the complete meaning of the original sentence. Figure 3 provides an example of PropNet.
PropNet is purely white-box since it is completely interpretable and transparent. It is also human-like because it has a hierarchy network structure to save the meaning in a sentence. With respect to semantics, unlike embedding models which focus more on the relationships between words and their contextual information, PropNet focuses on the true happenings in the world as expressed by a sentence. It achieves this by analyzing the core components of propositions, particularly the eight dimensions related to verbs. The principles behind PropNet’s design are very natural, without specially crafted artificial features. The comparison of PropNet and other representation methods is summarized in Table 1.
@@ 2 Relative Work
Neural network embedding approaches represent words as vectors in a continuous vector space. Representative methods include Word2Vec [42], FastText [10; 35; 34], Glove [45], BERT [20], SBERT [47], SimCSE [22] and AoE [39]. They train neural networks on an extremely large corpus, using different Language Modeling (LM) tasks as the objective. Since embeddings are automatically computed by backpropagation, feature engineering is completely eliminated. In addition, the embeddings provide deep semantic understanding of the text, as they capture the semantic aspects of language due to LM tasks as the training goal.
Embedding methods are considered as black-box algorithms because the embedding vectors are generally unexplainable. This characteristic surreptitiously encodes social biases contained in the training corpus [11; 14; 13; 44], and also raises concerns about trust [48; 49], safety [8; 46] and other issues [9]. Numerous efforts have been made to improve interpretability [30; 40] and remove biases [26; 25], but these problems have not been fundamentally solved.
All the aforementioned methods are not human-like, as their representation formats are either strings or vectors. Cognitive-Inspired Syntactic Models (CISM), such as Tree Adjoining Grammar [33], Combinatory Categorial Grammar [51], and Dependency Grammar [28], are based on principles of human language comprehension and production. These models use interpretable tree structures to represent sentences but primarily focus on syntactic organization. PropNet, which employs Dependency Grammar for parsing, enhances semantic representation through instance nodes and stamp nodes. Unlike CISM, PropNet uses a network structure rather than a tree structure to represent sentences. Table 1 summarizes the characteristics of these methods.
@@ 3 Methods
Figure 1: The framework of building PropNet. The type of the input sentence is calculated before splitting. After merging, all networks are integrated into one network as the PropNet of the input sentence. Note that for types Prop0 and Prop1 the splitting and merging phases are omitted.
@@
Table 2: Splitting rules for a Prop2 sentence according to its clause types. The condition token.pos_ in (VERB,AUX) is also required for all rules. We omit it to improve the clarity of the table. Note that token.pos_ and token.dep_ are the POS tagging and dependency parsing results from spaCy. The examples follow the format: Input Sentence -> (Main Proposition, Subordinate Proposition). The bold text in an example is the extracted verb token for the subordinate clause. Note that identifier_ccomp, identifier_pcomp, xcomp_identifier and identifier_csubj serve as placeholders for the relevant clauses.
To build a PropNet of a sentence, we first split the sentence into propositions (Section 3.1). Each proposition corresponds to a verb in the sentence. Then propositions are parsed to extract core tokens with respect to dimensions Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, and Source (Section 3.2). After that, each proposition is represented by a hierarchical network (Section 3.3). At last, all networks are merged into one network as the PropNet of the input sentence (Section 3.4). Figure 1 illustrates the framework of constructing PropNet. We introduce a method for comparing two PropNets, thereby unlocking the potential of the resulting PropNets for STS tasks (Section 3.5).
@@
  • P3+ indicates the set of pairs that includes at least one Prop3 sentence.
@@ 3.1 Splitting
For Prop0 and Prop1, splitting is unnecessary. At this stage, the input sentence itself serves as the output proposition. For Prop2, we split the input sentence into two Prop1 propositions, a main proposition (main clause) and a subordinate proposition (subordinate clause). Table 2 presents the splitting rules for different clause types. These rules extract the verb token of the subordinate clause. Then the subordinate clause is generated by iterating over the verb token’s subtree provided by spaCy parsing tool  [^1] For more details, see the spaCy documentation: https://spacy.io/usage/linguistic-features/#pos-tagging. For instance, in the Prop2 sentence “She thinks this is a good idea”, the splitting rules identify “is” as the verb token for the subordinate clause. The token “is” owns a subtree “this is a good idea” by dependency parsing.
The generated subordinate clause is then replaced by a clause identifier in the input sentence for five clause types Clausal Complement, Prepositional Complement, Open Clausal Complement, Subject Clause and Adverbial Clause. Their identifiers are notated as identifier_ccomp, identifier_pcomp, xcomp_identifier, identifier_csubj and identifier_advcl, respectively. For other clause types, the subordinate clause is replaced by an empty string. When the replacement is completed, the rest of the input sentence becomes the main proposition. Table 2 presents examples for each clause type. Note that identifier_advcl is applied only when the clause is preceded by the word “to” that indicates the purpose of an action. Figure 2 provides an example for this specific situation.
Clause identifiers are crucial for merging proposition representations into a unified network. In the above example, the main proposition is “She thinks identifier_ccomp” where identifier_ccomp would point to “this is a good idea” in the merging phase.
If the subordinate clause is non-finite, the subject of the main clause is spliced to it, forming a complete subordinate proposition. For instance, in “she wants to dance”, the subordinate proposition “she dance” uses the subject of the main proposition “she wants xcomp_identifier”. If the subordinate clause is finite, no subject insertion is required since the clause already contains its own subject. The splitting rules for relative clause type are more intricate and are detailed in Appendix A.2.
Situation becomes significantly more complex for Prop3+. We employ a backtracking algorithm to decompose the input sentence into Prop1 propositions by constructing a binary tree, as shown in Figure 2. The splitting method for Prop2 is invoked during each splitting step to extract a subordinate proposition Prop1, serving as the leaf node of the binary tree. Then the leaf node invokes the representation component to construct a hierarchical network, which is explained in Section 3.3. The leaf nodes indicate the order of the propositions contained in the long input sentence from left to right. In Figure 2, the input sentence has an ordered proposition list [“A black and white dog is jumping in the air to identifier_advcl”, “dog catch a Frisbee”, “a man is getting identifier_ccomp”, “his boat clean”, “he wants a water voyage”] according to the leaf nodes from left to right. This order is necessary when we compare two Prop3+ sentences for similarity, which is explained in Section 3.5.
@@ 3.2 Parsing
This phase extracts key tokens from a proposition with respect to specific dimensions. The intuition of designing these dimensions is that we require they can describe the meaning of a single action that happens in the real world, which is the key to understand the meaning of a complex sentence consisting of a series of actions. Therefore, eight critical dimensions related to the verb of a proposition are considered: Action, Subject, Object, Where, Auxiliary_Object, Goal, Reason, Source. The extracted words under each dimension are either describing a physical entity like “car” or an abstractive concept like “brand”. For these words, two additional dimensions Attribute and Part_of are considered for the descriptions of them. Explanations for these dimensions and extracting primary rules  [^2] Because of the variety of language, even a Prop1 sentence can have an extremely flexible structure. Therefore, minor rules are handcrafted for extraction which can be seen in the codes. with examples are listed in Table 3. The prepositions used by primary rules in the table are omitted for brevity. See Appendix A.3 for a detailed list of these prepositions. Extraction is implemented by resorting to spaCy part-of-speech tagging and dependency parsing tools  [^3] For more details, see the spaCy documentation: https://spacy.io/usage/linguistic-features/#pos-tagging.
@@ 3.3 Representing
As brain’s long-term memory stores information as a hierarchical network [18; 7; 37], PropNet represents a proposition using a six-level hierarchical network. Figure 3 (b) provides an example of how the sentence “A young man is playing an instrument in the garden” is represented. The network comprises three types of nodes: evolutionary nodes, developmental nodes and instance nodes which are represented as red, orange and green colors, respectively. In designing the network, we incorporate ideas similar to Nature Design and Nurture Beliefs , which are outlined in [52]. Evolutionary nodes correspond to Nature Design, which reflects human nature and is derived from heredity. They are denoted with a prefix “#”, like #action. Each dimension in the parsing phase has a corresponding evolutionary node. Developmental nodes correspond to Nurture Beliefs, which are learned from experience, and are named with a prefix “_”, such as _young. Each word in a sentence corresponds to a developmental node.
An evolutionary node cannot directly establish a connection with a developmental node. This connection is fulfilled by an instance node. Each evolutionary node points to at least one instance node. Each instance node can point to only one developmental node. They have the following three types:
@@
Figure 3 (a) provides a more lucid depiction of the six-level hierarchical structure of PropNet. An action instance node functions as the root and triggers #action. It also points to a developmental node _play. This action instance node is unique because there is only one verb in a proposition. Edges from #action connect evolutionary nodes #subject, #object, #where, #aux_obj, #goal, #source, and #reason, each of which corresponds to a parsing dimension described in Section 3.2. These nodes will establish connections with instance nodes that are created for the words parsed from each dimension. For example, #subject points to ins_entity_17, which has a link with _man. At level 4, these instance nodes activate two evolutionary nodes, #attr and #part_of, which align with the parsing dimensions Attribute and Part_of respectively. Similarly, #attr and #part_of establish links with instance nodes created for the words parsed under Attribute and Part_of. For instance, ins_entity_17 points to #attr and #attr points to _young.
@@
If during the parsing phase, no words are associated with a dimension, its corresponding evolutionary node will invoke an instance node that points to a default developmental node, _unknown. This ensures the network remains complete even when some information is missing. In Figure 3 (b), only #subject, #object and #where have valid instance nodes, which point to _man, _instrument and _garden, correspondingly. All other evolutionary nodes pointed to by #action are connected to instance nodes that point to _unknown, which is marked by a purple box in the figure.
@@ 3.5.1 Pair Type P1-
The comparison dimensions are formed by the Cartesian product of {#subject, #object, #where, #aux_obj, #goal, #source, #reason} and {$1$, #attr, #part_of}, further extended by {#action, #action|#subject|#where, #action|#object|#where}. For example, the dimension #action|#subject|#attr represents the attribute of an entity or concept that serves as the subject within the action. The reason of using the names of evolutionary nodes is this way of naming can reveals the route to locate the corresponding entity or concept instance nodes within the graph of PropNet. The explanations of all these 24 dimensions are provided in Appendix B.1.
By comparing the developmental nodes indicated by the dimensional instance nodes, we obtain a similarity code for each dimension. Concatenating these codes in sequence produces a difference vector $\mathbf{v_{1}}$. The similarity code primarily takes three values: 0 for identical, 1 for similar, 2 for different. A large language model (LLM) acts as the knowledge base to compute the similarity degree between two words. This strategy is word-level based, ignoring context-dependent meanings, e.g. polysemous words. Identification of the gloss that a word uses given a context is left in future work. Details on LLM choice and its prompts are explained in Appendix B.2. Additional similarity codes, and examples for each code are provided in Appendix B.3.
@@ 3.5.3 Pair Type P3+
As a Prop3+ sentence can contain multiple propositions, we restrict our consideration to its first four propositions in the order of the leaf nodes in its splitting binary tree from left to right. Given a P3+ pair, suppose the first sentence $s_{1}$ and the second sentence $s_{2}$ have ordered proposition lists $L_{1}=[p_{1},p_{2},p_{3},p_{4}]$ and $L_{2}=[q_{1},q_{2},q_{3},q_{4}]$ respectively. Note that $p_{2},p_{3},p_{4}$ from $s_{1}$ and $q_{4}$ from $s_{2}$ might be empty strings. We calculate the Proposition Overlap Degree (POD) as the number of overlapping words between two propositions. $L_{2}$ is aligned with $L_{1}$ according to POD by the following steps:
  1. (1)
    The PODs between $p_{1}$ and each one of $L_{2}$ are calculated. Suppose $q_{1}^{(1)}$ from $L_{2}$ has the largest POD which is greater than a threshold value 1, then $q_{1}^{(1)}$ is moved to the front of $L_{2}$. Other elements in $L_{2}$ keep their orders invariant. After this rearrangement, $L_{2}$ is notated as $L_{2}=[q_{1}^{(1)},q_{2}^{(1)},q_{3}^{(1)},q_{4}^{(1)}]$. If the largest POD is no greater than 1, then $L_{2}$ keeps its original order, which means $L_{2}=[q_{1}^{(1)},q_{2}^{(1)},q_{3}^{(1)},q_{4}^{(1)}]=[q_{1},q_{2},q_{3},q_{4}]$.
  2. (2)
    Suppose $L_{2}^{(1)}=[q_{2}^{(1)},q_{3}^{(1)},q_{4}^{(1)}]$. Compute the PODs between $p_{2}$ and each one of $L_{2}^{(1)}$. Similarly, check if the largest POD is greater than 1, and rearrange $L_{2}^{(1)}$ to $[q_{2}^{(2)},q_{3}^{(2)},q_{4}^{(2)}]$.
  3. (3)
    Suppose $L_{2}^{(2)}=[q_{3}^{(2)},q_{4}^{(2)}]$. Compute the PODs between $p_{3}$ and each one of $L_{2}^{(2)}$. Check if the largest POD is greater than 1, and rearrange $L_{2}^{(2)}$ to $[q_{3}^{(3)},q_{4}^{(3)}]$. Finally, $s_{2}$ has a new ordered proposition list $[q_{1}^{(1)},q_{2}^{(2)},q_{3}^{(3)},q_{4}^{(3)}]$ which is aligned to $s_{1}$ according to POD.
After alignment, there are four P1- pairs: $p_{1}$ and $q_{1}^{(1)}$, $p_{2}$ and $q_{2}^{(2)}$, $p_{3}$ and $q_{3}^{(3)}$, $p_{4}$ and $q_{4}^{(3)}$. We compute a difference vector $\mathbf{v_{1}}$ for each of them using previous computing method for a P1- pair. The final difference vector for this P3+ pair is the concatenation of these four $\mathbf{v_{1}}$s. Therefore, the difference vector $\mathbf{v_{3}}$ for P3+ has a size that is twice the length of $\mathbf{v_{2}}$ for P2. If two propositions are both empty, $\mathbf{v_{1}}$ is replaced by a vector of the same size, where all elements are 0. If one proposition is empty and the other one is not, $\mathbf{v_{1}}$ is replaced by a vector of the same size, where all elements are 2.
Let’s exemplify the above processes, considering the pair: “Three men are jumping off a wall” and “Three young men run, jump, and kick off of a Coke machine”. The second sentence consists of three propositions: “three young men run”, “three young men jump”, and “three young men kick off of a Coke machine”. After the alignment process, its proposition list is reordered to “three young men jump”, “three young men run”, “three young men kick off of a Coke machine”. The first sentence is compared with “three young men jump”, resulting a difference vector $\mathbf{v_{1}}$. Since “three young men run” and “three young men kick off of a Coke machine” have no counterparts in the first sentence, we concatenate $\mathbf{v_{1}}$ with two vectors of the same length as $\mathbf{v_{1}}$, padded with 2 to signify the two sentences are different in this part. Finally, the resulting vector is padded with 0 until it reaches the required length, indicating the two sentences are the same in this default part.